# HealthQ: Unveiling Questioning Capabilities of LLM Chains in Healthcare Conversations

### v3

@@ 5.1.2 Judgement Errors
5.2 Multi-shot Conversation with Memory
Real clinical interactions involve multiple QA rounds, which our current study does not fully capture due to text length and computational limitations. Future work with longer context windows and richer patient data can address this. As conversation complexity increases, it may be modeled as a dynamical system to study more intricate behaviors.
6 Conclusion
This paper presents a novel framework for evaluating the questioning capabilities of LLM healthcare chains, addressing a significant gap in the current research landscape. By implementing various state-of-the-art LLM chains and utilizing traditional NLP metrics for validation, we provide a comprehensive assessment of their ability to generate relevant and informative questions during patient interactions. Our approach includes the development of custom datasets and an evaluation framework that separates patient known knowledge from doctor questions, enabling precise evaluation. The results indicate that well-formulated questions enhance patient information gathering, thereby improving diagnostic accuracy and healthcare outcomes. Future work will focus on expanding the framework to support multi-shot conversations and further reducing biases in evaluation metrics. This framework sets a benchmark for future studies and aims to advance the effectiveness of LLM healthcare chains in real-world applications.

### v4

@@ V-A2 Judgement Errors
V-B Evaluation Limitations
A key limitation of our current framework is the reliance on LLMs for evaluation. Although we demonstrate consistency across different models (GPT-3.5, GPT-4, and Claude), this approach may overlook aspects of question quality that only human healthcare professionals could assess. Future work should incorporate practicing clinicians to evaluate generated questions and test them in real clinical settings. Such human-in-the-loop evaluation would provide insights into practical utility and effectiveness in actual medical conversations. Additionally, longitudinal studies with healthcare providers could help establish how these automated systems complement human medical expertise, particularly in assessing question appropriateness across specialties and alignment with diagnostic protocols.
VI Conclusion
This paper introduced HealthQ, a novel framework for evaluating the questioning capabilities of LLM healthcare chains, addressing the critical but underexplored role of question formulation in digital healthcare. By implementing advanced LLM chains, such as RAG and CoT, and leveraging an LLM judge alongside traditional NLP metrics like ROUGE and NER-based evaluations, we demonstrated the importance of well-crafted questions in eliciting informative patient responses. Validated across multiple LLM judge models and datasets, our results showed consistent performance, highlighting the framework’s robustness and potential for model-agnostic evaluation. HealthQ provides a comprehensive methodology for linking question quality to answer informativeness, setting a benchmark for advancing the effectiveness of LLMs in patient care. Future work will focus on expanding to multi-turn interactions, refining evaluation metrics, and exploring broader healthcare applications.